OpenStudy (anonymous):

a man has dimes and quarters worth 9.50. if he has 50 coins in all, how many dimes and how many quarters does he have ?

OpenStudy (tad1):

10d + 25 q + 9.50 d + q = 50 solve the system

OpenStudy (anonymous):

vierd

OpenStudy (tad1):

Solve for d, d= 50 - q Substitue for d in first equation, etc.

OpenStudy (anonymous):

can you show me how to solve the equation? im a little lost

OpenStudy (tad1):

sure .10d + .25q =9.50 = 10( 50 - q) + 25 q=9.50 5.00 - 10q +25q = 9,50 15q = 9.5-5 15 q = 4.00 q = 4.00/15 I don't think that works. I'll do it on paper and be right back.

OpenStudy (tad1):

Here we go. .10d + .25 q = 9.50 make the substitution .10( 50 - q) + .25q = 9.50 5 - .10q + .35q = 9.50 -,10q +.35 q = 4.50 .25 q = 4.50 q = 17 so d = 50 - 17 =

OpenStudy (anonymous):

33 dimes and 17 quarters, thanks !!